OpenStudy (pokiedokie):

Identify the conic section, any lines of symmetry, and the domain and range. x = 2y^2 + 5 , please help!!

OpenStudy (owlcoffee):

A conic section have a general equation of \[Ax^2+Bxy+Cy^2+Dx+Ey+F=0\] And this can be accomodated to fit any type of equation you have learned so far, this being: (1) A circle, (2) a parabola, (3) an elipse, (4) a hyperbola, (5) a product of lines. This is mostly evident through the discriminant \(\Delta=B^2-4AC\): (a) if \(\Delta>0\) then the conic section has hiperbolic genre (b) if \(\Delta=0\) then the conic section has Parabolic genre (c) if \(\Delta < 0\) then the conic section has eliptic genre. This is saved for when the conic section is complete and hard to represent, but in any case, your equation: \[x=2y^2+5\] Respects the form \(x=ay^2+by+c\) and that equation represented what geometric body?

OpenStudy (pokiedokie):

a hyperbola, and the domain would be 5, and i'm really not sure about anything else like the range? would it be 2?

OpenStudy (pokiedokie):

@Owlcoffee

OpenStudy (pokiedokie):

nvm i've figured some of it out, it's a parabola

OpenStudy (owlcoffee):

Good, now, what are the components of a parabola?

OpenStudy (pokiedokie):

the x and y intercepts and axis of symmetry?